On March 31, 2018, Chow Brothers, Inc., bought 5% of KT Manufacturing’s capital stock for $55.7 million. KT’s net income for the year ended December 31, 2018, was $81.9 million. The fair value of the shares held by Chow was $38.8 million at December 31, 2018. KT did not declare or pay a dividend during 2018.

Required:

1. Prepare all appropriate journal entries related to the investment during 2018.

- Record the purchase of KT Manufacturing’s capital stock for $55.7 million.

2. Assume that Chow sold the stock on January 20, 2019 for $31.9 million. Prepare the journal entries to record the sale.

- Record the entry for KT's net income.

3. Record any necessary entry to adjust the fair value on December 31, 2018.

4. Record any necessary entry to adjust the fair value on the date of sale.

5. Record the sale of stock on January 20, 2019 for $31.9 million.
Answer #1

1. Journal Entry in the books of Chow Brother's on Dec 31,2018

Investment Account Debit 55.7m

To Cash Account 55.7m

(Being the investment bought for the above value)

2. Journal Entry for the sale of the shares as on Jan 20,2019

Cash Account Debit 31.9m

Loss from the sale of investment Account Debit 23.8m

To Investment Account 55.7m

(Being investment sold with the loss)

3.Journal Entry for KT's Net Income

Profit and Loss Account Debit 81.9m

To KT's Capital Account 81.9m

(Being the profit transfered to Owner's Equity)

4.Calculation of the Fair value of the investment

5% of the Capital =55.7m

There total share Capital = 5% of the Capital*100/5

55.7m*100/5 =1114M

Total Profit for the year 81.9m

% of profit on Capital = Profit*100/Total Capital

81.9*100/1114 =7.35%

Profit Chow Brothers can expect from there investment for the year = 55.7*7.35/100

=4.09 m

Hence the stock bought on March 31 the profit will be = Number of months after investment/Total months in a year*4.09

=3.07

Hence the Fair Value as of Dec 31,2018 will be 55.7+3.07= 58.77

Journal entry to record the fair value in the Books of Chow

Investment Account Debit 3.07m

To Increase in the value of Investment 3.07m

(Being the Fair value of the investment recorded as of Dec 31)